# Vertex F-algebras and their associated Lie algebra

### v1

@@ 4.2 Meromorphicity
$i_{z,w}Y(a,F(z,w))\circ\mathcal{D}(w)b$ $=\mathcal{D}(w)\circ Y(a,z)b.$ (4.12)
@@
Putting $c=\mathbb{1}$ into (4.9) and using (4.11) shows
$F(z,w)^{N}\bigl{(}\mathcal{D}(w)\circ Y(a,z)b\bigr{)}=F(z,w)^{N}\bigl{(}Y(a,F(z,w))\circ\mathcal{D}(w)b\bigr{)}$ (4.13)
for some $N\geqslant 0.$ Since $\mathcal{D}(w)\circ Y(a,z)b\in V\llparenthesis z,w\rrparenthesis$ by (4.1) and (4.3) and $Y(a,F(z,w))\mathcal{D}(w)b\in V\llparenthesis z,w\rrparenthesis$ by (4.6) and (4.3), we can embed equation (4.13) into $V\llparenthesis z\rrparenthesis\llparenthesis w\rrparenthesis,$ where $F(z,w)$ is invertible. Equation (4.12) follows. ∎
@@
For each $a,b,c\in V,$ both sides of (4.9) are a Laurent series $p_{a,b,c}(z,w)\in V\llparenthesis z,w\rrparenthesis.$ The formal fraction
$f_{a,b,c}(z,w)=\frac{p_{a,b,c}(z,w)}{F(z,w)^{N}}\in V\llparenthesis z,w\rrparenthesis[F(z,w)^{-1}]$ (4.14)
is independent of the choice of $N.$ Under the homomorphisms $i_{w,z}$ and $i_{z,w}$ from $V\llparenthesis z,w\rrparenthesis[F(z,w)^{-1}]$ into $V\llparenthesis w\rrparenthesis\llparenthesis z\rrparenthesis$ and $V\llparenthesis w\rrparenthesis\llparenthesis z\rrparenthesis,$ respectively, we have
@@
$Y(Y(a,z)b,w)c=\sum\nolimits_{m,n\in\mathbb{Z}}\bigl{(}a_{(n)}(b)\bigr{)}_{(m)}(c)z^{-n-1}w^{-m-1}$
contains only finitely many negative powers of $z.$ Moreover, for a fixed power of $z,$ there are only finitely many negative powers of $w.$ Hence the left hand side of (4.9) belongs to $V\llparenthesis w\rrparenthesis\llparenthesis z\rrparenthesis.$ By definition of the substitution (3.13), the right hand side of (4.9) belongs to $V\llparenthesis z\rrparenthesis\llparenthesis w\rrparenthesis.$ Hence both sides of (4.9) are equal to a common series $p_{a,b,c}(z,w)\in V\llparenthesis z,w\rrparenthesis$ in the intersection. If we embed $p_{a,b,c}(z,w)$ into $V\llparenthesis w\rrparenthesis\llparenthesis z\rrparenthesis,$ we get $F(z,w)^{N}Y(Y(a,z)b,w)c$ and the series $F(z,w)$ is invertible in $V\llparenthesis w\rrparenthesis\llparenthesis z\rrparenthesis,$ so we may rearrange and get (4.15). Similarly, we may embed into $V\llparenthesis z\rrparenthesis\llparenthesis w\rrparenthesis$ and rearrange to get (4.16). ∎
@@
By (4.15) we can reexpress this identity as
$i_{z,w}f_{a,b,c}(z,w)=i_{z,w}f_{b,a,c}(\iota z,F(z,w)).$ (4.18)
in the ring $V\llparenthesis z\rrparenthesis\llparenthesis w\rrparenthesis.$ Perform the substitution $z\to F(v,\iota w)$ from (3.14) to get
@@
Since $f_{b,a,c}(F(w,\iota v),v)=w^{-N}p_{b,a,c}(F(w,\iota v),v)$ by (4.14) and $p_{b,a,c}(u,v)$ contains only finitely many negative powers of $u,$ there exists $M\geqslant 0$ such that $F(w,\iota v)^{M}f_{b,a,c}(F(w,\iota v),v)$ lies in $R\llparenthesis v,w\rrparenthesis.$ From (3.5) we thus have
$F(w,\iota v)^{M}i_{v,w}f_{b,a,c}(F(w,\iota v),v)=F(w,\iota v)^{M}i_{w,v}f_{b,a,c}(F(w,\iota v),v)$
@@
For all $a,b$ in a vertex $F$-algebra
$i_{z_{1},z_{0}}z_{2}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{0})}{z_{2}}\right)Y(Y(a,z_{0})b,z_{2})\\ =i_{z_{1},z_{2}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)Y(a,z_{1})Y(b,z_{2})\\ -i_{z_{2},z_{1}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)Y(b,z_{2})Y(a,z_{1}).$ (4.22)
@@
Let $c\in V.$ In the notation of (4.14), set $\phi_{a,b,c}(z_{0},z_{1},z_{2})=\frac{p_{a,b,c}(z_{0},z_{2})}{z_{1}^{N}}.$ Thus $\phi_{a,b,c}(z_{0},z_{1},z_{2})=f_{a,b,c}(z_{0},z_{2})$ if $F(z_{0},z_{2})=z_{1}.$ We have
$i_{z_{1},z_{2}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)\phi_{a,b,c}(z_{0},z_{1},z_{2})$ (4.23)
${\enskip=\enskip}$ $i_{z_{1},z_{2}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)f_{a,b,c}(F(z_{1},\iota z_{2}),z_{2})$
${\enskip=\enskip}$ $i_{z_{1},z_{2}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)Y(a,z_{1})Y(b,z_{2}).$
@@
$i_{z_{2},z_{1}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)\phi_{a,b,c}(z_{0},z_{1},z_{2})\\ =i_{z_{2},z_{1}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)Y(b,z_{2})Y(a,z_{1}).$ (4.24)
@@
$i_{z_{1},z_{0}}z_{2}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{0})}{z_{2}}\right)\phi_{a,b,c}(z_{0},z_{1},z_{2})$ (4.25)
${\enskip=\enskip}$ $i_{z_{1},z_{0}}z_{2}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{0})}{z_{2}}\right)f_{a,b,c}(z_{0},F(z_{1},\iota z_{0}))$
@@ 4.3 Construction of Lie bracket
$[a,b]=\operatorname{Res}_{z=0}^{F}Y(a,z)b\,dz.$ (4.26)
defines a Lie bracket on the quotient $V/\sum_{n\geqslant 1}\mathcal{D}^{(n)}(V).$
@@
$\operatorname{Res}^{F}_{z=0}Y(\mathcal{D}^{(m)}(a),z)b\,dz=\frac{1}{m!}\left.\frac{d^{m}}{dw^{m}}\right|_{w=0}\!\!\!\!\!\!\operatorname{Res}_{z=0}^{F}Y(\mathcal{D}(w)(a),z)b\,dz=0$ (4.27)
@@
$\operatorname{Res}_{z=0}\biggl{(}\frac{Y(a,z)\mathcal{D}^{(m)}(b)}{F^{1,0}(0,z)}\biggr{)}dz\in\sum_{n\geqslant 1}\mathcal{D}^{(n)}(V),$
@@
$Y(a,z)\circ\mathcal{D}(w)$ $=\mathcal{D}(w)\circ Y(a,F(z,\iota(w)))$
$=\mathcal{D}(w)\circ Y(\mathcal{D}(\iota(w))a,z)$
$=\sum\nolimits_{k,\ell\geqslant 0}\mathcal{D}^{(k)}\circ Y(\mathcal{D}^{(\ell)}(a),z)w^{k}\iota(w)^{\ell},$
which expresses $Y(a,z)\mathcal{D}^{(m)}(b)$ for $m\geq 1$ as a sum of terms with $k\geqslant 1$ or with $\ell\geqslant 1.$ The terms with $k\geqslant 1$ belong to $\sum_{n\geqslant 1}\mathcal{D}^{(n)}(V),$ while the terms with $\ell\geqslant 1$ are covered by (4.27).
@@
$\operatorname{Res}_{z=0}^{F}Y(\mathcal{D}(w)a,z)b\,dz$ $=\operatorname{Res}_{z=0}^{F}Y(a,F(z,w))b\,dz$
$=\sum_{n\geqslant 0}\bigl{(}\operatorname{Res}_{z=0}^{F}\mathcal{D}_{n}Y(a,z)b\bigr{)}w^{n}dz$
$=\operatorname{Res}_{z=0}^{F}Y(a,z)b\,dz$
@@
Putting these calculations into the formula (3.43) for interated $F$-residues with $(z_{0},z_{1},z_{2})\to(z,v,w)$ and using (3.31) yields a version of the Jacobi identity
$[a,[b,c]]-[b,[a,c]]=-[[b,a],c]$ (4.28)
Assuming a trivial center (meaning that $[d,c]=0$ for all $c$ implies $d=0$) this Jacobi identity implies also the skew symmetry of the Lie bracket. In general, (4.1) and (4.10) imply
$Y(a,z)b+\sum_{n\geqslant 1}\mathcal{D}^{(n)}(V)=Y(b,\iota(z))a+\sum_{n\geqslant 1}\mathcal{D}^{(n)}(V)$
@@
4.4 Example: Heisenberg Vertex F-Algebra
@@
$0\longrightarrow R\longrightarrow H_{F}\longrightarrow R\llparenthesis t\rrparenthesis\longrightarrow 0$ (4.29)
of the commutative Lie algebra $R\llparenthesis t\rrparenthesis$ by the cocycle
$c(f,g)=\operatorname{Res}^{F}_{z=0}\mathcal{D}_{1}^{F}(f)\cdot g=\operatorname{Res}_{z=0}f^{\prime}g.$ (4.30)
@@
$[b_{n},b_{m}]=n\delta_{n,-m}$ (4.31)
so we have obtained the ordinary Heisenberg Lie algebra except for having replaced $\mathbb{C}$ by the ring $R.$ Define the space of states
$V=R[b_{-1},b_{-2},\cdots].$
Write $\mathbb{1}\in V$ for the constant series $1,$ the vacuum. Define the vertex operator
$b(z)=\sum_{n\in\mathbb{Z}}b_{n}z^{-n-1}$
@@
$b_{n}\colon V\longrightarrow V,\quad b_{n}=\begin{cases}b_{n}\cdot(-)&(n<0),\\ n\frac{\partial}{\partial b_{-n}}&(n\geqslant 0).\end{cases}$
@@
$\mathcal{D}(z)=\sum_{n\geqslant 0}\mathcal{D}^{(n)}z^{n},\qquad\mathcal{D}^{(n)}\colon V\longrightarrow V,$
satisfying $\mathcal{D}(z)\mathbb{1}=\mathbb{1}$ and
$[\mathcal{D}(z),b(w)]=b(F(z,w)).$
These conditions actually determine $\mathcal{D}(z)$ completely. Written in components,
$[\mathcal{D}^{(n)},b_{m}]=\text{coefficient of $w^{m}$ in }(\mathcal{D}^{(n)}_{F}b)(w).$
This can be applied inductively to a state $b_{j_{1}}\cdots b_{j_{k}}$ until we reach $\mathcal{D}^{(n)}\mathbb{1}=0.$

### v2

@@ 4.2 Meromorphicity
$i_{z,w}Y(a,F(z,w))\circ\mathcal{S}(w)b$ $=\mathcal{S}(w)\circ Y(a,z)b.$ (4.8)
@@
Putting $c=\mathbbm{1}$ into (4.5) and using (4.7) shows
$F(z,w)^{N}\bigl(\mathcal{S}(w)\circ Y(a,z)b\bigr)=F(z,w)^{N}\bigl(i_{z,w}Y(a,F(z,w))\circ\mathcal{S}(w)b\bigr)$ (4.9)
for some $N\geqslant 0$. Since $\mathcal{S}(w)\circ Y(a,z)b\in V(\!(z,w)\!)$ by (4.1) and (4.2) and $Y(a,F(z,w))\mathcal{S}(w)b\in V(\!(z,w)\!)$ by (4.4) and (4.2), we can embed equation (4.9) into V((z))((w))V(\!(z)\!)(\!(w)\!), where $F(z,w)$ is invertible. Equation (4.8) follows. ∎
@@
For each $a,b,c\in V$, both sides of (4.5) are a Laurent series $p_{a,b,c}(z,w)\in V(\!(z,w)\!)$. The formal fraction
$f_{a,b,c}(z,w)=\frac{p_{a,b,c}(z,w)}{F(z,w)^{N}}\in V(\!(z,w)\!)\big[F(z,w)^{-1}\big]$ (4.10)
is independent of $N$. Under the homomorphisms $i_{w,z}$ and $i_{z,w}$ from $V(\!(z,w)\!)\big[F(z,w)^{-1}\big]$ into V((w))((z))V(\!(w)\!)(\!(z)\!) and V((w))((z))V(\!(w)\!)(\!(z)\!), respectively, we have
@@
$Y(Y(a,z)b,w)c=\sum_{m,n\in\mathbb{Z}}\bigl(a_{(n)}(b)\bigr)_{(m)}(c)z^{-n-1}w^{-m-1}$
contains only finitely many negative powers of $z$. Moreover, for a fixed power of $z$, there are only finitely many negative powers of $w$. Hence the left-hand side of (4.5) belongs to V((w))((z))V(\!(w)\!)(\!(z)\!). By definition of the substitution (3.11), the right-hand side of (4.5) belongs to V((z))((w))V(\!(z)\!)(\!(w)\!). Hence both sides of (4.5) are equal to a common series $p_{a,b,c}(z,w)\in V(\!(z,w)\!)$ in the intersection. If we embed $p_{a,b,c}(z,w)$ into V((w))((z))V(\!(w)\!)(\!(z)\!), we get $F(z,w)^{N}Y(Y(a,z)b,w)c$ and the series $F(z,w)$ is invertible in V((w))((z))V(\!(w)\!)(\!(z)\!), so we may rearrange and get (4.11). Similarly, we may embed into V((z))((w)){V(\!(z)\!)(\!(w)\!)} and rearrange to get (4.12). ∎
@@
By (4.11), we can reexpress this identity as
$i_{z,w}f_{a,b,c}(z,w)=i_{z,w}f_{b,a,c}(\iota z,F(z,w))$ (4.14)
in the ring V((z))((w))V(\!(z)\!)(\!(w)\!). Perform the substitution $z\to F(v,\iota w)$ from (3.12) to get
@@
Since $f_{b,a,c}(F(w,\iota v),v)=w^{-N}p_{b,a,c}(F(w,\iota v),v)$ by (4.10) and $p_{b,a,c}(u,v)$ contains only finitely many negative powers of $u$, there exists $M\geqslant 0$ such that $F(w,\iota v)^{M}f_{b,a,c}(F(w,\iota v),v)$ lies in R((v,w)){R(\!(v,w)\!)}. From (3.3) we thus have
$F(w,\iota v)^{M}i_{v,w}f_{b,a,c}(F(w,\iota v),v)=F(w,\iota v)^{M}i_{w,v}f_{b,a,c}(F(w,\iota v),v),$
@@
For all $a$, $b$ in a vertex $F$-algebra,
$i_{z_{1},z_{0}}z_{2}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{0})}{z_{2}}\right)Y(Y(a,z_{0})b,z_{2})$
$\qquad=i_{z_{1},z_{2}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)Y(a,z_{1})Y(b,z_{2})$
$\phantom{\qquad=}{}-i_{z_{2},z_{1}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)Y(b,z_{2})Y(a,z_{1}).$
@@
Let $c\in V$. In the notation of (4.10), set
$\phi_{a,b,c}(z_{0},z_{1},z_{2})=\frac{p_{a,b,c}(z_{0},z_{2})}{z_{1}^{N}}.$
Thus $\phi_{a,b,c}(z_{0},z_{1},z_{2})=f_{a,b,c}(z_{0},z_{2})$ if $F(z_{0},z_{2})=z_{1}$. We have
$i_{z_{1},z_{2}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)\phi_{a,b,c}(z_{0},z_{1},z_{2})$
[…]
@@
$i_{z_{2},z_{1}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)\phi_{a,b,c}(z_{0},z_{1},z_{2})$
$\qquad=i_{z_{2},z_{1}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)Y(b,z_{2})Y(a,z_{1}).$
@@
$i_{z_{1},z_{0}}z_{2}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{0})}{z_{2}}\right)\phi_{a,b,c}(z_{0},z_{1},z_{2})$
@@
$\qquad=i_{z_{1},z_{0}}z_{2}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{0})}{z_{2}}\right)Y(Y(a,z_{0})b,F(z_{1},\iota z_{0}))c.$ (4.19)
@@ 4.3 Construction of Lie bracket
$a,b]=\operatorname{Res}_{z=0}^{F}Y(a,z)b{\rm d}z$ (4.20)
defines a Lie bracket on the quotient $V/\sum_{n\geqslant 1}\mathcal{S}^{(n)}(V)$.
@@
$\operatorname{Res}^{F}_{z=0}Y\bigl(\mathcal{S}^{(m)}(a),z\bigr)b{\rm d}z=\frac{1}{m!}\left.\frac{d^{m}}{{\rm d}w^{m}}\right|_{w=0}\operatorname{Res}_{z=0}^{F}Y(\mathcal{S}(w)(a),z)b{\rm d}z=0$ (4.21)
@@
$\operatorname{Res}_{z=0}\biggl(\frac{Y(a,z)\mathcal{S}^{(m)}(b)}{F^{1,0}(0,z)}\biggr){\rm d}z\in\sum_{n\geqslant 1}\mathcal{S}^{(n)}(V),$
@@
$Y(a,z)\circ\mathcal{S}(w)$ $=\mathcal{S}(w)\circ Y(a,F(z,\iota(w)))=\mathcal{S}(w)\circ Y(\mathcal{S}(\iota(w))a,z)$
$=\sum_{k,\ell\geqslant 0}\mathcal{S}^{(k)}\circ Y\big(\mathcal{S}^{(\ell)}(a),z\big)w^{k}\iota(w)^{\ell},$
which expresses $Y(a,z)\mathcal{S}^{(m)}(b)$ for $m\geqslant 1$ as a sum of terms with $k\geqslant 1$ or with $\ell\geqslant 1$. The terms with $k\geqslant 1$ belong to $\sum_{n\geqslant 1}\mathcal{S}^{(n)}(V)$, while the terms with $\ell\geqslant 1$ are covered by (4.21).
@@
$\operatorname{Res}_{z=0}^{F}Y(\mathcal{S}(w)a,z)b{\rm d}z$ $=\operatorname{Res}_{z=0}^{F}Y(a,F(z,w))b{\rm d}z=\sum_{n\geqslant 0}\bigl(\operatorname{Res}_{z=0}^{F}\mathcal{S}_{n}Y(a,z)b\bigr)w^{n}{\rm d}z$
$=\operatorname{Res}_{z=0}^{F}Y(a,z)b{\rm d}z,$
@@
Putting these calculations into the formula (3.29) for iterated $F$-residues with $(z_{0},z_{1},z_{2})\to(z,v,w)$ and using (3.22) yields a version of the Jacobi identity
[a,[b,c]][b,[a,c]]=[[b,a],c].[a,[b,c]]-[b,[a,c]]=-[[b,a],c].
Assuming a trivial center (meaning that $[d,c]=0$ for all $c$ implies $d=0$), this Jacobi identity implies also the skew symmetry of the Lie bracket. In general, (4.1) and (4.6) imply
$Y(a,z)b+\sum_{n\geqslant 1}\mathcal{S}^{(n)}(V)=Y(b,\iota(z))a+\sum_{n\geqslant 1}\mathcal{S}^{(n)}(V)$
@@
4.4 Example: Heisenberg vertex $\boldsymbol{F}$-algebra
@@
$0\longrightarrow R\longrightarrow H_{F}\longrightarrow R(\!(t)\!)\longrightarrow 0$
of the commutative Lie algebra R((t))R(\!(t)\!) by the cocycle
$c(f,g)=\operatorname{Res}^{F}_{z=0}\mathcal{S}_{1}^{F}(f)\cdot g=\operatorname{Res}_{z=0}f^{\prime}g.$
@@
$b_{n},b_{m}]=n\delta_{n,-m},$ (4.22)
so we have obtained the ordinary Heisenberg Lie algebra except for having replaced $\mathbb{C}$ by the ring $R$. Define the space of states $V=R[b_{-1},b_{-2},\dots]$. Write $\mathbbm{1}\in V$ for the constant series $1$, the vacuum. Define the vertex operator
$b(z)=\sum_{n\in\mathbb{Z}}b_{n}z^{-n-1},$
@@
$b_{n}\colon\ V\longrightarrow V,\qquad b_{n}=\begin{cases}b_{n}\cdot(-)&(n<0),\\ n\dfrac{\partial}{\partial b_{-n}}&(n\geqslant 0).\end{cases}$
@@
$\mathcal{S}(z)=\sum_{n\geqslant 0}\mathcal{S}^{(n)}z^{n},\qquad\mathcal{S}^{(n)}\colon\ V\longrightarrow V,$
satisfying $\mathcal{S}(z)\mathbbm{1}=\mathbbm{1}$ and $\mathcal{S}(z),b(w)]=b(F(z,w))$. These conditions actually determine $\mathcal{S}(z)$ completely. Written in components,
$\big[\mathcal{S}^{(n)},b_{m}\big]=\text{coefficient of $w^{m}$ in }\bigl(\mathcal{S}^{(n)}_{F}b\bigr)(w).$
This can be applied inductively to a state $b_{j_{1}}\cdots b_{j_{k}}$ until we reach $\mathcal{S}^{(n)}\mathbbm{1}=0$.
Acknowledgements
The author would like to thank an anonymous referee for several suggestions, in particular for the referee’s comments on Remark 4.4.